OpenStudy (anonymous):

This table shows the mass in grams m of the radioactive substance iodine-131 remaining in a container t days after the beginning of an experiment.

OpenStudy (anonymous):

Write a function that models the data.

OpenStudy (mathmale):

This problem is a bit more challenging than the 3 before it in that you are not told what kind of model to use.

OpenStudy (mathmale):

You could try a linear model (which would probably be a poor fit to the data), a quadratic model or an exponential model.

OpenStudy (anonymous):

Could I use a quadratic? Or would I use something else?

OpenStudy (anonymous):

I could try doing a quadratic.

OpenStudy (mathmale):

Let me summarize those 3: y=ax + b y=ax^2 + bx + c (our old friend) y=ae^(bx) exponential model

OpenStudy (anonymous):

I had some help on this question from a friend, he told me to use y=ae^(bx), but I didn't know how to use it.

OpenStudy (mathmale):

Remember, we don't always get Perfect Fits when we develop models. Its very possible you could create an acceptable expo model or an acceptable quad model.

OpenStudy (mathmale):

I believe I've seen this question before and that the other person and I came up with an exponential model. Basically, you're doing the same thing, Ricky: finding suitable values for the constants a, or a and b, or a and b and c.

OpenStudy (mathmale):

In this case, with 2 constants undetermined, we'd need only 2 points from the table to determine those 2.

OpenStudy (mathmale):

Please go ahead and pick 2 pts from the table. Typel them here as (x,y).

OpenStudy (anonymous):

(1,917.40), (2,841.62)

OpenStudy (anonymous):

Then I substitute for x and y right?

OpenStudy (mathmale):

Fine. Now stick those values into y=ae^(bx). for exampel, from the first point, 841.62 = a e^(b*2), where x = 2 and y =841.62. To answer your qeustion, yes.

OpenStudy (mathmale):

Why don't you use your point (2, 841.62) but use ( 0 , 1000) for the 2nd? You'll see why in a moment.

OpenStudy (mathmale):

Using these 2 points, please come up with 2 equations in a and b.

OpenStudy (anonymous):

i can do that!

OpenStudy (mathmale):

Much like you did earlier when trying to find a, b and c for a quadratic model.

OpenStudy (mathmale):

sounds wonderful, Ricky: "I can do that!" more power to you.

OpenStudy (anonymous):

so this is what I have so far: 1000 = ae^(b(0)) -> ae = 1000 841.62 = ae^(b(2)) -> ae^(2b) = 841.62

OpenStudy (mathmale):

The first one is actually ae^0 = 1000. Note that e^0 is 1, whereas e =2.718 approx. Big difference.

OpenStudy (mathmale):

So,. solve 1000 ae^0 for a.

OpenStudy (mathmale):

The value of a is 1000. Then yo ur expo model becomes y=1000e^(bx) Substitute the data from the other point: x=2 and y=841.62. This is all the info you need to solve for b.

OpenStudy (anonymous):

And if that is right (which it is), then i could do this to solve for a: 1a = 1000 -> 1000 Then I can do this to solve for b: (1000)e^(2b(2)) = 841.62 correct?

OpenStudy (mathmale):

Exactly. Think: what's the easiest way to isolate b?

OpenStudy (mathmale):

think of properties of log and expo functions.

OpenStudy (anonymous):

Well first you would have to multiply 2b and 2 together to get 4b. Then after that, would you divide?

OpenStudy (mathmale):

Are you u sing properties of logs and expo functions here?

OpenStudy (anonymous):

Oh, no. I wasn't

OpenStudy (mathmale):

841.62=1000e^(2b) your best bet is to take the natural log (ln) of boths ides. ln 841.62 = ln 1000 + ln e^(2b) Hope this looks familiar for you. ln 841.62-ln 1000 = 2b. Before we do any more be certain that what I've demo'd here is clear for y ou. If it isn't we'll need to talk.

OpenStudy (anonymous):

Yes, it is clear!

OpenStudy (mathmale):

ln 841.62-ln 1000 = 6.735 - 8.294 = 2 b

OpenStudy (mathmale):

I'd forgotten that you don't have a calculator. How are you computing your logs?

OpenStudy (anonymous):

b = 417.306

OpenStudy (anonymous):

I don't have graphing calculator like yours, but I do have one that does log functions.

OpenStudy (mathmale):

6.735 - 8.294 = 2 b => all you have to do is to divide both sides of this equation by 2, and you'll have your b. Note that b is qute different from 417.3.

OpenStudy (mathmale):

try it, please.

OpenStudy (mathmale):

I come up with the exponential model y=1000 e ^ (-0.0862 x). To check this, simply pick another point from the table of x and y values, substitute them into this model and determine whether the resulting equation is true or false.

OpenStudy (anonymous):

I got: -0.7795 = b

OpenStudy (mathmale):

You are doing the right thing. Why your b and my b are different, I don't know. I was just beginning to check my work.

OpenStudy (mathmale):

While I do that plese think of any last min. questions you'd liek to ask. I have to excuse myself very soon as I am dri8ving towards Los Angeles this morning for an 11:30 a.m. apopiontment.

OpenStudy (anonymous):

okay! Let me see.

OpenStudy (mathmale):

I tried the same work again, same approach, and still get b = -0.0862.

OpenStudy (anonymous):

hmm, okay. Well, that was the last one that I really needed help on! Thank you so much! I think I can take it from here. If i have a question, I will just post it on here for whenever.

OpenStudy (mathmale):

Try this If 841.62=1000 e^(bx), 841.62 ------ = e^(bx) 1000 Taking the log of both sides (after re-writing 841.62/1000 as 0.8416), ln 0.8416 = 2b b=ln 0.8416/2 = ?

OpenStudy (anonymous):

-0.865598

OpenStudy (mathmale):

I should have enclosed that log in parentheses for added clarity. b=(ln 0.8416)/2 = ?

OpenStudy (anonymous):

-0.0862252

OpenStudy (mathmale):

OK, I think we've spent enuf time on that one. Very happy to work with you! Probably see you not today but tomorrow or later this week.

OpenStudy (mathmale):

There you go. Correct result.

OpenStudy (anonymous):

Thank you! Have a great day!

OpenStudy (mathmale):

Then your expo model is y = 1000 e^(-0.0862x).

OpenStudy (mathmale):

Bye!